a steamer moving in a river takes m seconds to go 'a' meters downstream and n seconds to go 'a' meters upstream.Find the speed of the river water and the speed of the steamer relative to the river water in m/sec.

Grade:12

1 Answers

Shubham Singh
8 years ago
vm + um = a or m(v+u) = a
vn – un = a or n(v-u) = a
v = speed of steamer
u = speed of river
speed of steamer relative to water will be v – u = vr
form above 2 equations
m(vr + 2u) = a
n vr = a
from these 2 u can calculate u and vr